# Spacetime Quasicrystals

### v1

@@ II.3 Symmetries of Lattices
We define a root of the lattice $\Lambda$ to be a vector $\bf r$} of positive norm $N(r)=\langle{\bf r},{\bf r}\rangle>0$ for which the associated reflection is a symmetry of $\Lambda$. If $\Lambda$ is self-dual, its roots are just the lattice vectors ${\bf r}\in\Lambda$ whose norm $N({\bf r})=\langle r,r\rangle$ equals $1$ or $2$. In particular, the lattice vectors with $N(r)=1$ or $N(r)=2$ are called the short roots and long roots, respectively.
@@ II.5 Algebraic Number Theory
A monic polynomial $P(x)$ is a polynomial in $x$ whose leading coeffcient is $1$. To any element $\alpha$ in the algebraic number field $K$ is associated a minimal polynomial $P_{\alpha,\mathbb{Q}}(x)$, which is the lowest-order monic polynomial with coefficients in $\mathbb{Q}$, having $\alpha$ as a root. The roots of $P_{\alpha,\mathbb{Q}}(x)$ are called the algebraic conjugates (or just conjugates) of $\alpha$. [^4] If the extension $K/\mathbb{Q}$ is normal, one can equivalently say the conjugates of $\alpha$ are the images of $\alpha$ under the field automorphisms of $K$ that fix the elements of $\mathbb{Q}$ – a key idea in Galois theory.
@@ III.1.1 Construction
  • Choice 3: a subset $\cal W$} of $V_{{\rm in}}$ (the “window”) or a function $W:V_{{\rm in}}\to\mathbb{C}$ (the “weighting function).
@@
$W({\bf v}_{{\rm in}})=\left\{\begin{array}[]{ll}1&{\bf v}_{{\rm in}}\in\mathcal{W},\\ 0&{\bf v}_{{\rm in}}\notin\mathcal{W}.\end{array}\right$. (25)
@@ III.1.3 Self-Dual Quasicrystals
First consider periodic crystals. By the Poisson summation formula, the distribution
@@
Also note that the window scheme, where $W$ is the indicator function (25) for some compact window $\mathcal{W}$, does not yield a self-dual quasicrystal. In particular, $W$ is non-vanishing on a compact region in $V_{{\rm in}}$, implying that $\Lambda_{{\rm ph}}^{{\bf t},W}$ has delta functions on a finite-density (non-dense) set of points in $V_{{\rm ph}}$; while $\widehat{W}$ is non-vanishing on a non-compact region of $V_{{\rm in}}$, implying that $\widehat{\Lambda}_{{\rm ph}}^{{\bf t},W}$ has delta functions on a dense set of points in $V_{{\rm ph}}$ (i.e. there are delta functions arbitrarily close to any point in $V_{{\rm ph}}$). By contrast, in a self-dual quasicrystal, $W$ and $\widehat{W}$ are both non-vanishing on a non-compact region of $V_{{\rm in}}$, implying that $\Lambda_{{\rm phys}}^{{\rm t},W}$ and $\widehat{\Lambda}_{{\rm phys}}^{\,{\rm t},W}$ both have delta functions on a dense set of points in $V_{{\rm ph}}$. For illustration, we present examples of both types (the 8-fold symmetric Ammann-Beenker tiling and its self-dual variant) in Sec. III.2.2 and Fig. 7.
@@ III.2.1 Construction
  • Choice 2: a noncrystallographic subgroup $G\subset{\rm Aut}(\Lambda)$, and a decomposition $V=V_{{\rm ph}}\oplus V_{{\rm in}}$ that is invariant under $G$: $\Pi_{{\rm ph}}=g\Pi_{{\rm ph}}g^{-1}$ (or equivalently $\Pi_{{\rm ph}},g]=0$), $\forall g\in G$.
@@ III.2.2 Example 1: Ammann-Beenker Tiling
Choice 1: As our lattice $\Lambda^{{\bf t}}$, we choose the standard integer lattice ($\Lambda=\mathbb{Z}^{4}$) with no translation (${\bf t}={\bf 0}$). ${\rm Aut}(\Lambda)$ is the Coxeter group $B_{4}$, whose fundamental roots $r_{i}$ are the columns of the matrix
@@
From each $r_{i}$ we construct the corresponding reflection matrix $R_{i}$ from Eq. (10), and hence the corresponding Coxeter element
@@ III.3.1 The torus parametrization
Let us describe three important relationships between quasicrsytals described by different values of the paramter $\bf t$}:
  • Torus parameterization [6]. First note that if we shift ${\bf t}\to{\bf t}^{\prime}={\bf t}+\bm{\lambda}$, with $\bm{\lambda}\in\Lambda$, the two quasicrystals are identical:
@@
  • Local indistinguishability. Two quasicrystals $\Lambda_{{\rm ph}}^{{\rm t},W}$ with same $\Lambda$ and $\cal W$} (or $W$), but different $\bf t$} are locally indistinguishable [8, 5] [^5] The term “locally indistinguishable” used here is synonymous with the earlier term “locally isomorphic” introduced in [47]. Informally, this means that any finite patch found somewhere in one quasicrystal can also be found somewhere in the other.
@@
    Quasicrystals $\Lambda_{{\rm ph}}^{{\rm t},W}$ with the same $\Lambda$ and $\cal W$} (or $W$), but different $t$, are in the same local isomorphism class, and the quasicrystals in this equivalence class are in one-to-one correspondence with the points on the torus ${\bf t}\in T^{d}=V/\Lambda$.
    Note that, for the same $\Lambda$ and $\cal W$} (or $W$), but different translation parameters $\bf t$} and $\bf t$$^{\prime}$, where $\bf t$$^{\prime}$ cannot be written as ${\bf t}+\Delta{\bf t}+\bm{\lambda}$ (with $\Delta{\bf t}\in V_{{\rm ph}}$ and $\bm{\lambda}\in\Lambda$), the two quasicrystals $\Lambda_{{\rm ph}}^{{\bf t},W}$ and $\Lambda_{{\rm ph}}^{{\bf t}^{\prime},W}$ are not equal up to translation: so they are locally isomorphic but globally distinct. The local isomorophism class contains an uncountable infinity of locally isomorophic yet globally distinct quasicrystals.
III.3.2 (Quasi)translational invariance
Since $V_{{\rm ph}}$ is totally irrational relative to $\Lambda$ (i.e. $V_{{\rm ph}}$ does not intersect any points in $\Lambda$, besides the origin), the resulting C&P pattern is aperiodic: i.e. it has no non-zero vector $\Delta{\bf v}_{{\rm ph}}\in V_{{\rm ph}}$ s.t. $\Lambda_{{\rm ph}}^{{\rm t},W}({\bf v}_{{\rm ph}})=\Lambda_{{\rm ph}}^{{\rm t},W}({\bf v}_{{\rm ph}}+\Delta{\bf v}_{{\rm ph}})$.
@@
III.3.3 (Quasi)orientational invariance
In the sCNP scheme (where the lattice $\Lambda$ has automorphism group ${\rm Aut}(\Lambda)$, while the projector $\Pi_{{\rm ph}}$ and the weighting function $W$ are both invariant under a subgroup $G\subset{\rm Aut}(\Lambda)$): for ${\bf t}={\bf 0}$, the resulting quasicrystal $\Lambda_{{\rm ph}}^{{\bf 0},W}$ also has exact orientational invariance under $G$:
@@
III.3.4 (Quasi)Scale invariance
@@ III.3.5 Local scale invariance
To apply this to sC&P , consider two patterns, $P=\Lambda_{{\rm ph}}^{{\bf t},{\cal W}}({\bf v}_{{\rm ph}})$ and $P^{\prime}=\Lambda_{{\rm ph}}^{{\rm t},\widetilde{{\cal W}}}({\bf v}_{{\rm ph}})$, with $\widetilde{{\cal W}}=\lambda_{{\rm in}}{\cal W}$.
@@
In other words, the quasicrystal $\Lambda_{{\rm ph}}^{{\bf t},{\cal W}}$ (obtained using the old translation $\bf t$} and the new window $\widetilde{{\cal W}}$) is the same as the quasicrystal $\lambda_{{\rm ph}}\Lambda_{{\rm ph}}^{\widetilde{{\bf t}},{\cal W}}$ (obtained by using the old window $\cal W$} and the new translation $\bf t$}, and then rescaling the resulting quasicrystal by $\lambda_{{\rm ph}}$). And this, in turn is locally isomorphic to the quasicrystal $\lambda_{{\rm ph}}\Lambda_{{\rm ph}}^{{\bf t},W}$ (by replacement $\widetilde{t}\to t$). Hence, $P$ is locally isomorphic to $\lambda_{{\rm ph}}P^{\prime}$.
@@
For example, in the AB tiling, the original window $\cal W$} is a regular octagon, while the rescaled window $\widetilde{{\cal W}}$ is a smaller regular octagon (scaled down by the factor $1+\sqrt{2}$). Since each of these shapes may be construct from a finite numer of unions and intersections of the other shape, it follows that the AB tiling has local scale invariance – the inflation/deflation rule illustrated in Fig. 8.
@@ III.3.6 Global scale invariance
In other words, Eq. (43) says that, as we rescale the physical space $V_{{\rm ph}}\to\lambda_{{\rm ph}}V_{{\rm ph}}$, the underlying dense subset $\Lambda_{{\rm ph}}\subset V_{{\rm ph}}$ is mapped to itself, one-to-one; and, moreover, as we map any point $x_{{\rm ph}}\in\Lambda_{{\rm ph}}$ to the new point $x_{{\rm ph}}^{\prime}=\lambda_{{\rm ph}}x_{{\rm ph}}\in\Lambda_{{\rm ph}}$, we find that the delta function peak living at the original point $x_{{\rm ph}}$ was higher (by a factor of $\lambda_{{\rm in}}^{k}$) than the delta function peak living at the new point $x_{{\rm ph}}^{\prime}$. We say such a quasicrystal has exact global scale invariance with weight $k$.
So far we have focused on the case ${\bf t}=0$, since it yields exact global scale invariance. But it is also singular since the delta function peaked at the origin (${\bf v}_{{\rm ph}}=0$) is multiplied by an infinite coefficient $W(0)=\infty$.
@@ IV.1.1 Lorentzian Lattices
In Sec. II.1, we introduced lattices living in Euclidean space $V=\mathbb{R}^{d}$. To generalize the discussion, we now replace $\mathbb{R}^{d}$ (with $d$ spatial dimensions) to $\mathbb{R}^{(s,t)}$ (with $s$ spatial and $t$ temporal dimensions). In other words, we replace the positive-definite inner product $\langle{\bf v},{\bf w}\rangle=v^{\alpha}\delta_{\alpha\beta}w^{\beta}$ on $\mathbb{R}^{d}$ by the non-degenerate but indefinite symmetric bilinear form $\langle{\bf v},{\bf w}\rangle=v^{\alpha}\eta_{\alpha\beta}w^{\beta}$ on $\mathbb{R}^{(s,t)}$ where, in an orthonormal (Cartesian) basis, $\eta_{\alpha\beta}$ is
@@
  • Odd self-dual lattices: For any indefinite $V=\mathbb{R}^{(s,t)}$ (with $s,t$ both $\geq 1$), there is a unique odd self-dual lattice (denoted $I_{s,t}$) which, in an orthonormal (Cartesian) basis, consists of all vectors $x\in\mathbb{R}^{(s,t)}$ with integer components $x^{\mu}$:
    $I_{s,t}=\{x|x^{\mu}\in\mathbb{Z}\;(\mu=1,\ldots,s+t)\}.$ (45a)
@@
  • Even self-dual lattices: For any indefinite $V=\mathbb{R}^{(s,t)}$ (with $s,t$ both $\geq 1$) whose signature $\sigma\equiv s-t$ is also a multiple of 8, we also have a unique even self-dual lattice (denoted ${\rm II}_{s,t}$) which, in an orthonormal (Cartesian) basis, consists of all vectors $x\in\mathbb{R}^{(s,t)}$ with (i) components $x^{\mu}$ all in $\mathbb{Z}$ or all in $\mathbb{Z}+\frac{1}{2}$; and (ii) even inner product with the vector $n=(1,\ldots,1$):
    ${\rm II}_{s,t}$ $=$ $\{x|\,{\it all}\;x^{\mu}\in\mathbb{Z}\;{\rm or}\,{\it all}\,x^{\mu}\in\mathbb{Z}+\frac{1}{2};$ (45b)
    $\quad{\rm and}\;\langle n,x\rangle\in 2\mathbb{Z}\;{\rm for}\;n=(1,\ldots,1)\}.$

### v2

@@ II.3 Symmetries of Lattices
A root of the lattice $\Lambda$ is a vector $\bf r$} of positive norm $N({\bf r})=\langle{\bf r},{\bf r}\rangle>0$ for which the associated reflection is a symmetry of $\Lambda$. If $\Lambda$ is self-dual, its roots are just the lattice vectors ${\bf r}\in\Lambda$ whose norm $N({\bf r})$ equals $1$ or $2$. In particular, the lattice vectors with $N({\bf r})=1$ or $N({\bf r})=2$ are called the short roots and long roots, respectively.
@@ II.5 Algebraic Number Theory
A monic polynomial $P(x)$ is a polynomial in $x$ whose leading coefficient is $1$. To any element $\alpha$ in the algebraic number field $K$ is associated a minimal polynomial $P_{\alpha,\mathbb{Q}}(x)$, which is the lowest-order monic polynomial with coefficients in $\mathbb{Q}$, having $\alpha$ as a root. The roots of $P_{\alpha,\mathbb{Q}}(x)$ are called the algebraic conjugates (or just conjugates) of $\alpha$. [^4] If the extension $K/\mathbb{Q}$ is normal, one can equivalently say the conjugates of $\alpha$ are the images of $\alpha$ under the field automorphisms of $K$ that fix the elements of $\mathbb{Q}$ – a key idea in Galois theory.
@@ III.1.1 Construction
  • Choice 3: a subset $\cal W$} of $V_{{\rm in}}$ (called the “window”) or a function $W:V_{{\rm in}}\to\mathbb{C}$ (called the “weighting function”).
@@
$W({\bf v}_{{\rm in}})=\left\{\begin{array}[]{ll}1,&\text{ if }{\bf v}_{{\rm in}}\in\mathcal{W},\\ 0,&\text{ if }{\bf v}_{{\rm in}}\notin\mathcal{W}.\end{array}\right$. (25)
@@ III.1.3 Self-Dual Quasicrystals
First, consider periodic crystals. By the Poisson summation formula, the distribution
@@
Also note that the window scheme, where $W$ is the indicator function (25) for some compact window $\mathcal{W}$, does not yield a self-dual quasicrystal. In particular, $W$ is non-vanishing on a compact region in $V_{{\rm in}}$, implying that $\Lambda_{{\rm ph}}^{{\bf t},W}$ has delta functions on a finite-density (non-dense) set of points in $V_{{\rm ph}}$; while $\widehat{W}$ is non-vanishing on a non-compact region of $V_{{\rm in}}$, implying that $\widehat{\Lambda}_{{\rm ph}}^{{\bf t},W}$ has delta functions on a dense set of points in $V_{{\rm ph}}$ (i.e. there are delta functions arbitrarily close to any point in $V_{{\rm ph}}$). By contrast, in a self-dual quasicrystal, $W$ and $\widehat{W}$ are both non-vanishing on a non-compact region of $V_{{\rm in}}$, implying that $\Lambda_{{\rm phys}}^{{\bf t},W}$ and $\widehat{\Lambda}_{{\rm phys}}^{\,{\bf t},W}$ both have delta functions on a dense set of points in $V_{{\rm ph}}$. For illustration, we present examples of both types (the 8-fold symmetric Ammann-Beenker tiling and its self-dual variant) in Sec. III.2.2 and Fig. 7.
@@ III.2.1 Construction
  • Choice 2: a non-crystallographic subgroup $G\subset{\rm Aut}(\Lambda)$, and a decomposition $V=V_{{\rm ph}}\oplus V_{{\rm in}}$ that is invariant under $G$: $\Pi_{{\rm ph}}=g\Pi_{{\rm ph}}g^{-1}$ (or equivalently $\Pi_{{\rm ph}},g]=0$), $\forall g\in G$.
@@ III.2.2 Example 1: Ammann-Beenker Tiling
Choice 1: As our lattice $\Lambda^{{\bf t}}$, we choose the standard integer lattice ($\Lambda=\mathbb{Z}^{4}$) with no translation (${\bf t}={\bf 0}$). ${\rm Aut}(\Lambda)$ is the Coxeter group $B_{4}$, whose fundamental roots ${\bf r}_{i}$ are the columns of the matrix
@@
From each ${\bf r}_{i}$ we construct the corresponding reflection matrix $R_{i}$ from Eq. (10), and hence the corresponding Coxeter element
@@ III.3.1 The Torus Parametrization
Let us describe three important relationships between quasicrystals described by different values of the parameter $\bf t$}:
  • Torus parametrization [6]. First note that if we shift ${\bf t}\to{\bf t}^{\prime}={\bf t}+\bm{\lambda}$, with $\bm{\lambda}\in\Lambda$, the two quasicrystals are identical:
@@
  • Local indistinguishability. Two quasicrystals $\Lambda_{{\rm ph}}^{{\bf t},W}$ with same $\Lambda$ and $\cal W$} (or $W$), but different $\bf t$} are locally indistinguishable [8, 5] [^5] The term “locally indistinguishable” used here is synonymous with the earlier term “locally isomorphic” introduced in [48]. Informally, this means that any finite patch found somewhere in one quasicrystal can also be found somewhere in the other.
@@
    Quasicrystals $\Lambda_{{\rm ph}}^{{\bf t},W}$ with the same $\Lambda$ and $\cal W$} (or $W$), but different $\bf t$}, are in the same local isomorphism class, and the quasicrystals in this equivalence class are in one-to-one correspondence with the points on the torus ${\bf t}\in T^{d}=V/\Lambda$.
    Note that, for the same $\Lambda$ and $\cal W$} (or $W$), but different translation parameters $\bf t$} and $\bf t$$^{\prime}$, where $\bf t$$^{\prime}$ cannot be written as ${\bf t}+\Delta{\bf t}+\bm{\lambda}$ (with $\Delta{\bf t}\in V_{{\rm ph}}$ and $\bm{\lambda}\in\Lambda$), the two quasicrystals $\Lambda_{{\rm ph}}^{{\bf t},W}$ and $\Lambda_{{\rm ph}}^{{\bf t}^{\prime},W}$ are not equal up to translation: so they are locally isomorphic but globally distinct. The local isomorphism class contains an uncountable infinity of locally isomorphic yet globally distinct quasicrystals.
III.3.2 (Quasi-)Translational Invariance
Since $V_{{\rm ph}}$ is totally irrational relative to $\Lambda$ (i.e. $V_{{\rm ph}}$ does not intersect any points in $\Lambda$, besides the origin), the resulting C&P pattern is aperiodic: i.e. it has no non-zero vector $\Delta{\bf v}_{{\rm ph}}\in V_{{\rm ph}}$ s.t. $\Lambda_{{\rm ph}}^{{\bf t},W}({\bf v}_{{\rm ph}})=\Lambda_{{\rm ph}}^{{\bf t},W}({\bf v}_{{\rm ph}}+\Delta{\bf v}_{{\rm ph}})$.
@@
III.3.3 (Quasi-)Orientational Invariance
In the sC&P scheme (where the lattice $\Lambda$ has automorphism group ${\rm Aut}(\Lambda)$, while the projector $\Pi_{{\rm ph}}$ and the weighting function $W$ are both invariant under a subgroup $G\subset{\rm Aut}(\Lambda)$): for ${\bf t}={\bf 0}$, the resulting quasicrystal $\Lambda_{{\rm ph}}^{{\bf 0},W}$ also has exact orientational invariance under $G$:
@@
III.3.4 (Quasi-)Scale Invariance
@@ III.3.5 Local Scale Invariance
To apply this to sC&P , consider two patterns, $P=\Lambda_{{\rm ph}}^{{\bf t},{\cal W}}({\bf v}_{{\rm ph}})$ and $P^{\prime}=\Lambda_{{\rm ph}}^{{\bf t},\widetilde{{\cal W}}}({\bf v}_{{\rm ph}})$, with $\widetilde{{\cal W}}=\lambda_{{\rm in}}{\cal W}$.
@@
In other words, the quasicrystal $\Lambda_{{\rm ph}}^{{\bf t},{\cal W}}$ (obtained using the old translation $\bf t$} and the new window $\widetilde{{\cal W}}$) is the same as the quasicrystal $\lambda_{{\rm ph}}\Lambda_{{\rm ph}}^{\widetilde{{\bf t}},{\cal W}}$ (obtained by using the old window $\cal W$} and the new translation $\bf t$}, and then rescaling the resulting quasicrystal by $\lambda_{{\rm ph}}$). And this, in turn is locally isomorphic to the quasicrystal $\lambda_{{\rm ph}}\Lambda_{{\rm ph}}^{{\bf t},W}$ (by replacement $\widetilde{\bf t}\to{\bf t}$). Hence, $P$ is locally isomorphic to $\lambda_{{\rm ph}}P^{\prime}$.
@@
For example, in the AB tiling, the original window $\cal W$} is a regular octagon, while the rescaled window $\widetilde{{\cal W}}$ is a smaller regular octagon (scaled down by the factor $1+\sqrt{2}$). Since each of these shapes may be construct from a finite number of unions and intersections of the other shape, it follows that the AB tiling has local scale invariance – the inflation/deflation rule illustrated in Fig. 8.
@@ III.3.6 Global Scale Invariance
In other words, Eq. (43) says that, as we rescale the physical space $V_{{\rm ph}}\to\lambda_{{\rm ph}}V_{{\rm ph}}$, the underlying dense subset $\Lambda_{{\rm ph}}\subset V_{{\rm ph}}$ is mapped to itself, one-to-one; and, moreover, as we map any point ${\bf x}_{{\rm ph}}\in\Lambda_{{\rm ph}}$ to the new point ${\bf x}_{{\rm ph}}^{\prime}=\lambda_{{\rm ph}}{\bf x}_{{\rm ph}}\in\Lambda_{{\rm ph}}$, we find that the delta function peak living at the original point ${\bf x}_{{\rm ph}}$ was higher (by a factor of $\lambda_{{\rm in}}^{k}$) than the delta function peak living at the new point ${\bf x}_{{\rm ph}}^{\prime}$. We say such a quasicrystal has exact global scale invariance with weight $k$.
So far we have focused on the case ${\bf t}=0$, since it yields exact global scale invariance. But it is also singular since the delta function peaked at the origin (${\bf v}_{{\rm ph}}=0$) is multiplied by an infinite coefficient $W({\bf 0})=\infty$.
@@ IV.1.1 Lorentzian Lattices
In Sec. II.1, we introduced lattices living in Euclidean space $V=\mathbb{R}^{d}$. To generalize the discussion, we now replace $\mathbb{R}^{d}$ (with $d$ spatial dimensions) to $\mathbb{R}^{s,t}$ (with $s$ spatial and $t$ temporal dimensions). In other words, we replace the positive-definite inner product $\langle{\bf v},{\bf w}\rangle=v^{\alpha}\delta_{\alpha\beta}w^{\beta}$ on $\mathbb{R}^{d}$ by the non-degenerate but indefinite symmetric bilinear form $\langle{\bf v},{\bf w}\rangle=v^{\alpha}\eta_{\alpha\beta}w^{\beta}$ on $\mathbb{R}^{s,t}$ where, in an orthonormal (Cartesian) basis, $\eta_{\alpha\beta}$ is
@@
  • Odd self-dual lattices: for any indefinite $V=\mathbb{R}^{s,t}$ (with $s,t$ both $\geq 1$), there is a unique odd self-dual lattice (denoted $I_{s,t}$) which, in an orthonormal (Cartesian) basis, consists of all vectors ${\bf x}\in\mathbb{R}^{s,t}$ with integer components ${x}^{\mu}$:
    $I_{s,t}=\{{\bf x}|x^{\mu}\in\mathbb{Z}\;(\mu=1,\ldots,s+t)\}.$ (45a)
@@
  • Even self-dual lattices: for any indefinite $V=\mathbb{R}^{s,t}$ (with $s,t$ both $\geq 1$) whose signature $\sigma\equiv s-t$ is also a multiple of 8, we also have a unique even self-dual lattice (denoted ${\rm II}_{s,t}$) which, in an orthonormal (Cartesian) basis, consists of all vectors ${\bf x}\in\mathbb{R}^{s,t}$ with (i) components $x^{\mu}$ all in $\mathbb{Z}$ or all in $\mathbb{Z}+\frac{1}{2}$; and (ii) even inner product with the vector $n=(1,\ldots,1$):
    ${\rm II}_{s,t}$ $=$ $\{{\bf x}|\,{\it all}\;x^{\mu}\in\mathbb{Z}\;{\rm or}\,{\it all}\,x^{\mu}\in\mathbb{Z}+\frac{1}{2};$ (45b)
    $\quad{\rm and}\;\langle{\bf n},{\bf x}\rangle\in 2\mathbb{Z}\;{\rm for}\;{\bf n}=(1,\ldots,1)\}.$